Question 18

2 Points

Humpback whales can dive to 176 m depth where the pressure is 1.78 MPa. Assuming that the whales body temperature is constant (36.40 oC), what volume will the whales lungs now have? 2,048.7636 L (Assume that the pressure at the surface is 1.000 atm or 101.3 kPa and the initial volume of the lungs is 3.600 x 103 L, give your answer to 3 significant figures, do not use scientific notation)

  • BLANK-1
Answer #1

Hi,

Hope you are doing well.


__________________________________________________________________________

According to Boyle's law, we know that, at constant temperature, pressure is inversely proportional to the volume of the gas.

Here, given that, temperature is constant.

Initial pressure, Pi = 101.3 kPa = 101300 Pa

Final pressure, Pi= 1.78 x 106 Pa

Initial volume, V = 3600 L

Therefore final volume is given by,

V_{f}=\frac{P_{i}V_{i}}{P_{f}}=\frac{3600\;L\times 101300\;Pa}{1.78\times10^{6}\;Pa}

..V = 204.88 L

Rounding to 3 significant figures,

:: Ve= 205 L


__________________________________________________________________________

Hope this helped for your studies. Keep learning. Have a good day.

Feel free to clear any doubts at the comment section.


Please don't forget to give a thumps up.

Thank you. :)